Divide rs 600 in amount x,y,z in the ratio 1:2:3?
Answer by TimothyLamb(4379)

---
x = (1/6)*600 = 100
y = (2/6)*600 = 200
z = (3/6)*600 = 300
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php